OpenStudy (anonymous):

Write the standard form of the equation of each line given the slope and y-intercept. slope = 3/5 y-int = 5

OpenStudy (anonymous):

standard form is y=mx+b

OpenStudy (anonymous):

So that would be 5=mx+b But there isn't an x variable.

OpenStudy (texaschic101):

standard form is Ax + By = C

OpenStudy (anonymous):

the slope is m

OpenStudy (texaschic101):

first put the equation is y = mx + b form,,,,m is your slope and b is your y intercept. So slope intercept form is : y = 3/5x + 5 But we need it in standard form of Ax + By = C y = 3/5x + 5 --- multiply the entire equation by the common denominator 5, because standard form will not have fractions. 5y = 3x + 25 --- subtract 3x from both sides -3x + 5y = 25 --- multiply by -1 to make x positive 3x - 5y = -25 .<=== standard form

OpenStudy (anonymous):

Thank you very much. Although I'm pretty sure you're not supposed to give out answers :p Thank you for explaining it though.

OpenStudy (anonymous):

im sorry i had a blonde momnent and switched the two ...